N2(8) + 3F2(g) → 2 NF3() AH 298 = - 264 kJ mol-; AS98 = - 278 J K-mol- The following questions relate to the synthesis rea
Answer #1

a. ΔG0298 = ΔH0298 - T* ΔS0298

   ΔH0298 = -264 kJ / mole , T = 298 K and ΔS0298 = -278 J / K mole = -0.278 kJ / K mole

ΔG0298 = -264 - 298*-0.278 = -181.156 kJ / mole

b.  ΔG0rxn = -RT lnKeq

Here, Keq = 1 lnKeq = ln 1 = 0 -RT lnKeq = 0   ΔG0rxn = -RT lnKeq = 0

So,  ΔG0rxn = ΔH0 - T* ΔS0 = 0

-264 - T*-0.278 = 0 T = 264/0.278 = 949.64 K

c. Standard enthalpy change for the formation of 2 moles of NF3 = -264 kJ / mole

So, Standard enthalpy change for the formation of 1 moles of NF3 = -264/2 kJ / mole = -132 kJ / mole

So, Standard enthalpy change for the formation of 0.256 moles of NF3 = -132*0.256 = - 33.792 kJ  

d.Initially there are N2 and F2 only. So, one molecule of NF3 will form 3 new N-F bonds. So, 2 molecule of NF3 will form 2*3 = 6 new N-F bonds

e. ΔH0rxn = E(bonds breaking ) - E(bond-forming)

-264 kJ/ mole  = E (N≡N) + 3* E(F-F) - 2*3* E(N-F)

-264 kJ / mole = 946 kJ / mole + 3* E(F-F) - 6*272 kJ / mole

So, 3* E(F-F) = -264 +6*272 -946 = 422 kJ  

  E(F-F) = 422/3 = 140.66 kJ / mole